a seedman has seeds worth $.70 per pound and seeds worth $.90 per pound. How many pounds of each must he use to make 300 pounds worth $.75 per pound?
Answer by checkley77(12844)

.90x+.70(300-x)=.75*300
.90x+210-.70x=225
,20x=225-210
.20x=15
x=15/.20
x=75 pounds of $.90 seeds are used.
300-75=225 pounds of $.70 seeds are used.
Proof:
.90*75+.70*225=225
67.5+157.5=225
225=225